OpenStudy (adrimit):

does anyone know a way to solve a system of five linearly independent equations in five variables?

OpenStudy (mww):

Easiest way is to use a matrix of 5x5 and find its reduced row echelon form, or row echelon form and back substitute.

OpenStudy (adrimit):

I found a way to solve a system of linear equations using multidimensional spheres

OpenStudy (wolf1728):

WOW- I found a way to that with 2, 3 and even 4 unknowns but 5 is a bit much. I used Cramer's Rule to do that but even using that for even 3 unknowns gets a little cumbersome.

OpenStudy (adrimit):

my method involves finding the center of a multidimensional sphere given a set of points and their tangent displacements. the center is the solution of the system of equations. notice that it is very cumbersome to calculate the solution of a set of linear equations for any more variables than three using traditional methods. using my algorithm, the equaltions solve themselves

OpenStudy (wolf1728):

So, you do have a way to solve for 5 equations in 5 unknowns?

OpenStudy (sachintha):

If your method is correct, please educate us. :)

OpenStudy (mww):

the question is whether this extends to n equations with n unknowns?

OpenStudy (adrimit):

first, I shall define the matrix representation of a sphere in n dimensions \[ m+1\;points\;are\;needed\;to\;define\;a\;sphere\;of\;m\;dimensions \\x_{(m+1)k}:k^{th}\;component\;of\;m^{th}\;point\;(given) \\d_m^2:tangent\;displacement\;of\;the\;m^{th}\;point\;(given) \\x_{0k}:k^{th}\;component\;of\;center\;(to\;be\;determined) \\R^2:radius\;squared\;(to\;be\;determined) \\\sum^{n}_{k=1}\left(x_{(m+1)k}-x_{0k}\right)^2=R^2+d_m^2 \\\sum^{n}_{k=1}\left(x_{(m+1)k}^2-2x_{(m+1)k}x_{0k}+x^2_{0k}\right)=R^2+d^2_m \] now, subtract the equation with m=0 from all the others the result is: \[\sum^n_{k=1}\left(\left(x^2_{(m+1)k}-x^2_{1k}\right)-2\left(x_{(m+1)k}-x_{1k}\right)x_{0k}\right)=d^2_m-d_0^2 \\this\;is\;linear\;in\;the\;components\;of\;the\;center \\so\;a\;matrix\;can\;be\;formed\;to\;solve\;for\;the\;center\;in\;terms\;of\;the \\given\;points\;and\;displacements \] however at this point, this is no simpler than the direct method of solving the original system of equations. Fortunately, the spheres can be used to build the solution geometrically I will post the result soon

ganeshie8 (ganeshie8):

just commenting to follow this

OpenStudy (adrimit):

Thank you for your interest. the final equation in my previous post is linear in the coordinates of the center, so I can rewrite it in a more familiar form: \[\sum^{n}_{k=1}2\left(x_{(m+1)k}-x_{1k}\right)x_{0k}=\sum^{n}_{k=1}\left(x^2_{(m+1)k}-x^2_{1k}\right)-\left(d^2_m-d^2_0\right) \\\;\\\;\\If\;A\;is\;the\;coefficient\;matrix,\;then\;the\;cell\;of\;A\;at\ \\the\;j^{th}\;row\;and\;the\;k^{th}\;column\;is: \\A(j,k)=2(x_{(j+1)k}-x_{1k}) \\\;\\\;\\If\;B\;denotes\;the\;constants\;vector,\;then\;the\;j^{th}\;component\;of\;B\;is:\; \\B(j)=\sum^{n}_{k=1}\left(x^2_{(j+1)k}-x^2_{1k}\right)-\left(d^2_j-d^2_0\right) \\\;\\\;\\The\;system\;of\;equations\;can\;be\;defined\;in\;tensor\;notation\;by\: \\AX=B \\\;\\\;\\Solve\;for\;X\;(center\;coordinates): \\X=A^{-1}B \] this is the way to find the center using matrices, there is a more elegant way to do it with geometry

OpenStudy (zzr0ck3r):

calculators

OpenStudy (adrimit):

computers

OpenStudy (zzr0ck3r):

I wonder how CAS's do it.

OpenStudy (owen3):

What do you mean by 'tangent displacement'.

OpenStudy (adrimit):



OpenStudy (adrimit):

the distance from the point where the tangent intersects the circle to the point itself

OpenStudy (owen3):

So the m+1 points are not on the actual sphere, they are off the sphere, or randomly placed, but the tangential displacements are given?

OpenStudy (inkyvoyd):

What's the O time of your solution?

OpenStudy (adrimit):

the points and their displacements are given. If the displacement is 0, then the point is on the surface of the sphere.

OpenStudy (adrimit):

linear using recursion

OpenStudy (owen3):

\[ \Large x_{o_k} , x_{1_k} , x_{2_k} .. x_{m_k} \]

OpenStudy (owen3):

Nevermind, we are solving for the center.

OpenStudy (adrimit):

bingo

OpenStudy (owen3):

\[ \Large x_{o_k} , x_{1_k} , x_{2_k} .. x_{m_k} , x_{m+1_k} \]

OpenStudy (adrimit):

\[x_{0_k}\;is\;the\;k^{th}\;component\;of\;the\;center\]

OpenStudy (owen3):

I have a feeling we might need a double sum.

OpenStudy (adrimit):

no, we need a system of equations

OpenStudy (owen3):

$$ \Large \sum_{i=1}^{m+1} \sum^{n}_{k=1}\left(x_{i_k}-x_{0_k}\right)^2=R^2+d_m^2 $$

OpenStudy (adrimit):

each point other than the one chosen as the origin has its own equation

OpenStudy (adrimit):

the term origin here is a little misleading because it does not actually need to be 0

OpenStudy (adrimit):

\[x_{1k}\;is\;the\;k^{th}\;element\;of\;the\;origin\]

OpenStudy (owen3):

Are you changing your problem above. you used \( \large x_{o_k} \) to denote the center. Sorry these subscripts are tripping me up. copy pasted: $$m+1\;points\;are\;needed\;to\;define\;a\;sphere\;of\;m\;dimensions \\x_{(m+1)k}:k^{th}\;component\;of\;m^{th}\;point\;(given) \\d_m^2:tangent\;displacement\;of\;the\;m^{th}\;point\;(given) \\x_{0k}:k^{th}\;component\;of\;center\;(to\;be\;determined) \\R^2:radius\;squared\;(to\;be\;determined) $$ Is this a correct edit, change m to m+1 $$x_{(m+1)k}:k^{th}\;component\;of\;m+1^{th}\;point\;(given) $$

OpenStudy (adrimit):

yeah that seems right